# PrivaDE: Privacy-preserving Data Evaluation for Blockchain-based Data Marketplaces

### v1

@@ 5.1. Dataset Scoring Algorithm
Given a representative set $D_{R}$, the parties compute a score using three criteria—diversity, uncertainty, and loss—where diversity and uncertainty follow standard active-learning practice.
Diversity. The subset should be diverse, meaning elements of $\mathcal{S}$ are sufficiently different on average to avoid redundancy. We denote diversity by $\mathcal{D}$, instantiated, e.g., as a statistic (mean, median) of pairwise distances. Diverse-set constructions are common in coreset and batch active learning (Sener and Savarese, 2018).
Uncertainty. Uncertainty indicates potential informativeness: high uncertainty suggests the model benefits from additional training on such samples. A typical measure is the entropy of $M$’s predictive distribution; margin-based variants are also used. We write the uncertainty functional as $\mathcal{U}$.
Loss. Relevance can be captured by the loss incurred by $M$ on Bob’s data (Yoo and Kweon, 2019). If $y^{\prime}=M_{\theta}(x)$, then $\ell(y^{\prime},y)$ quantifies the discrepancy between prediction and label; assuming valid data, larger loss indicates greater potential value for training.
@@
For multi-class image classification, choose cross-entropy loss $\ell_{\mathrm{CE}}$, margin-based uncertainty $u=\frac{1}{k}\sum_{i}(\max p_{i}-\text{second-max }p_{i})$ (Settles, 2009), diversity $d=$ mean pairwise $\|x-x^{\prime}\|$ over $D_{R}^{x}$, and $f(l,u,d)=\alpha l+\beta u+\gamma d$ for weights $\alpha,\beta,\gamma$.
@@ 5.2. The PrivaDE Protocol
    Verified inference. The model is evaluated on $D_{R}$, and the cut-and-choose protocol $\mathsf{CnCZK}$ is used to verify sampled layer computations.
@@
Formally, we prove that the protocol realizes $\mathcal{F}_{\mathsf{Score}}$, which securely computes Algorithm 1. In particular, Alice’s input (the model weights) and Bob’s input are protected except for some intended leakage. This leakage profile is fully specified by $\mathcal{F}_{\mathsf{Score}}$ in Fig. 5.
@@
Parameters: model $M=C\circ B\circ A$, loss $\ell$, uncertainty function $\mathcal{U}$, diversity function $\mathcal{D}$, scoring function $f$, representative set selection $\mathcal{R}$, distance threshold $d$, outlier ratio $\delta$, commitment parameter $\mathsf{pp}$ (1) $\mathcal{F}_{\mathsf{Score}}$ receives $I_{R}$ (indices of representative set) from $P_{2}$. (2) $\mathcal{F}_{\mathsf{Score}}$ receives a list of challenge indices II from $P_{1}$, and sends $I$ to $P_{2}$. (3) $\mathcal{F}_{\mathsf{Score}}$ receives $D^{x}_{B|I}$ (The set $D_{B}^{x}$ indiced by II) and $D^{x}_{R|I}$ (respective points in $D^{x}_{R}$ closest to each $D^{x}_{B|I}$) from $P_{2}$. For each $i\in I$ $(x_{i}\in D^{x}_{B|I})$, $\mathcal{F}_{\mathsf{Score}}$ checks that there exists some $x_{j}\in D_{R|I}^{x}$ such that $\|x_{i}-x_{j}\|<d$. If this check fails for at least $\delta\cdot|I|$ indices, the functionality sends $\mathsf{abort}$ to both $P_{1}$ and $P_{2}$. (4) $\mathcal{F}_{\mathsf{Score}}$ receives $\theta_{A}$ from $P_{1}$ and $D^{x}_{R}$ from $P_{2}$, then computes $\overrightarrow{a_{A}}\leftarrow A_{\theta_{A}}(D^{x}_{R})$ and sends $\overrightarrow{a_{A}}$ to $P_{2}$. (5) $\mathcal{F}_{\mathsf{Score}}$ receives $\theta_{B}$ from $P_{1}$, sends $\theta_{B}$ to $P_{2}$, then computes $\overrightarrow{a_{B}}\leftarrow B_{\theta_{B}}(\overrightarrow{a_{A}})$ and sends $\overrightarrow{a_{B}}$ to $P_{1}$ and $P_{2}$. (6) $\mathcal{F}_{\mathsf{Score}}$ receives $\theta_{C}$ from $P_{1}$, computes $\overrightarrow{y^{\prime}}\leftarrow C_{\theta_{C}}(\overrightarrow{a_{B}})$. (7) $\mathcal{F}_{\mathsf{Score}}$ receives $D^{x}_{B},D^{y}_{R}$ from $P_{2}$ and calculates the score: $l=\frac{1}{k}\sum_{i=1}^{k}\ell(y^{\prime}_{i},y_{i}),\quad u=\mathcal{U}(\{y^{\prime}_{1},\ldots,y^{\prime}_{k}\}),\quad d=\mathcal{D}(D_{B}^{x},D_{R}^{x})$, and finally $\phi\leftarrow f(l,u,d)$. It then sends $\phi$ to both $P_{1}$ and $P_{2}$.
Figure 5. Functionality $\mathcal{F}_{\mathsf{Score}}$ for computing the score of a dataset.
@@
We are now ready to formally describe the protocol PrivaDE that realizes $\mathcal{F}_{\mathsf{Score}}$ (Fig. 5). PrivaDE uses the following tools:
@@
  • $(\mathsf{SetupCom},\mathsf{Commit},\mathsf{Open})$ is a commitment scheme satisfying Definition B.1.
  • $(\mathsf{SetupZKP},\mathcal{P},\mathcal{V})$ is an NIZKPoK system satisfying Definition 2.2.3, defined for the NP language $L_{\textsf{ML-Inf}}$ (Sec. 4.4). We specify the protocol with end-to-end ZKPs, but in practice this can be replaced by the weaker $\mathsf{CnCZK}$ audit with tunable detection probability. Our proof-of-concept implementation uses EZKL (zkonduit, 2023).
@@
  • $\Pi_{\mathsf{Inference}}$ is a 2-party protocol satisfying Definition 2.2.4, which securely realizes $\mathcal{F}_{\mathsf{Inference}}$ (Fig. 6) with malicious security. We instantiate this with SPDZ2k (Cramer et al., 2018) in MP-SPDZ (Keller, 2020).
@@
The Setup phase. The setup phase is an offline stage before the protocol between Alice and Bob begins. In this phase:
@@
  • Alice and Bob communicate with a trusted third party to generate setup parameters $\mathsf{pp}\leftarrow\mathsf{SetupCom}(1^{\lambda})$ and $\mathtt{CRS}\leftarrow\mathsf{SetupZKP}(1^{\lambda})$.
@@
  • Bob verifies the authenticity of his dataset $D_{B}$ with a data authority, which then issues a signature attesting to the authenticity of $D_{B}$.
@@
  • Alice optionally performs model distillation locally (no verification required), then interacts with a model authority to split the model $M$ into $C\circ B\circ A$, with the authority validating that the DINA reconstruction SSIM is <0.3<0.3 (Sec. 4.2.2).
@@
This is the only phase that requires a trusted third party. We then proceed to the main protocol, described in Fig. 8.
@@
Parties: Model owner $P_{1}$, data owner $P_{2}$ Private parameters: Dataset $D_{B}$ (held by $P_{2}$); model weights $\theta_{A},\theta_{C}$ (held by $P_{1}$) Public parameters: Commitment parameter $\mathsf{pp}$, representative set size $k$, distance threshold $d$, outlier ratio $\delta$, model architecture $M$, number of challenges $m$, model weights $\theta_{B}$ Stage 0: Commitments (1) $P_{1}$ computes $(\mathtt{com}_{A},r_{A})\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{A})$ and $(\mathtt{com}_{C},r_{C})\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{C})$, and sends $\mathtt{com}_{A},\mathtt{com}_{C}$ to $P_{2}$. (2) For each $(x_{i},y_{i})\in D_{B}$, $P_{2}$ computes $(\mathtt{com}_{x_{i}},r_{x_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},x_{i})$ and $(\mathtt{com}_{y_{i}},r_{y_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},y_{i})$, and sends $\mathtt{com}_{x_{i}},\mathtt{com}_{y_{i}}$ to $P_{1}$. Stage 1: Representative Set Selection (3) $P_{2}$ runs the representative set selection algorithm $\mathcal{R}$ locally to select a subset $D_{R}$ of size $k$ from $D_{B}$. It then computes $d^{\prime}$ as the $(1-\delta)$-th percentile of the pairwise distances $\Big\{\min_{b\in D_{R}^{x}}\|a-b\|:a\in D_{B}^{x}\setminus D_{R}^{x}\Big\}$. If $d^{\prime}>d$, $P_{2}$ aborts the protocol. Otherwise, he sends indices $I_{R}$ to $P_{1}$. (4) $P_{1}$ and $P_{2}$ runs the challenge protocol $\mathsf{CP}$, where $P_{1}$ supplies a list of challenge indices II of size m. If the challenge protocol fails, $P_{1}$ aborts. Stage 2: Model Inference (5) $P_{1}$ and $P_{2}$ run the 2PC protocol $\Pi_{\mathsf{Inference}}$ to evaluate block AA on $D_{R}^{x}$. $P_{1}$ inputs $(\theta_{A},r_{A})$; $P_{2}$ inputs $D_{R}^{x}$ (and the corresponding opening randomness). $P_{2}$ receives the activations $\overrightarrow{a_{A}}=\{a_{A,i}\}_{i}$ and per-sample commitments $\{(\mathtt{com}_{a_{i}},r_{a_{i}})\}_{i}$; both parties obtain $\{\mathtt{com}_{a_{i}}\}_{i}$. (6) $P_{2}$ locally computes $\overrightarrow{a_{B}}\leftarrow B_{\theta_{B}}(\overrightarrow{a_{A}})$. He generates a ZK proof $\pi_{B}$ for the language $L_{\textsf{ML-Inf}}$ (public weights, hidden inputs) with witness $(\overrightarrow{a_{A}},\{r_{a_{i}}\}_{i})$, and sends $(\overrightarrow{a_{B}},\pi_{B})$ to $P_{1}$. (7) $P_{1}$ verifies $\pi_{B}$; if verification fails, she aborts. (8) $P_{1}$ computes $\overrightarrow{y^{\prime}}\leftarrow C_{\theta_{C}}(\overrightarrow{a_{B}})$ and for each $i$ computes $(\mathtt{com}_{y^{\prime}_{i}},r_{y^{\prime}_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},y^{\prime}_{i})$. She then produces a ZK proof $\pi_{C}$ for $L_{\textsf{ML-Inf}}$ (hidden weights, public inputs) with witness $(\theta_{C},r_{C},y^{\prime}_{1},\ldots,y^{\prime}_{k},r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}})$, and sends $\{\mathtt{com}_{y^{\prime}_{i}}\}_{i=1}^{k}$ and $\pi_{C}$ to $P_{2}$. (9) $P_{2}$ verifies $\pi_{C}$; if verification fails, he aborts. Stage 3: Dataset Scoring (10) $P_{1}$ and $P_{2}$ run the maliciously secure 2PC protocol $\Pi_{\mathsf{SubScore}}$. $P_{1}$ inputs $\big(y^{\prime}_{1},\ldots,y^{\prime}_{k},\,r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}},\,\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}\big)$; $P_{2}$ inputs $\big(D_{B}^{x},\,y_{1},\ldots,y_{k},\,r_{y_{1}},\ldots,r_{y_{k}},\,\mathtt{com}_{y_{1}},\ldots,\mathtt{com}_{y_{k}}\big)$. Both parties receive the score $\phi$..

### v2

@@ 4.1. Dataset Scoring Algorithm
As proven in numerous active-learning literature (e.g. (settles2009active; dasgupta2011two)), the use of active learning heuristics often gives good indicator of the overall dataset quality. Hence, given a representative set $D_{R}$, the parties compute a score using three criteria - diversity, uncertainty, and loss, where diversity and uncertainty follow standard active-learning practice.
Diversity. The subset should be diverse, with elements sufficiently different to avoid redundancy. We denote diversity by $\mathcal{D}$, instantiated as a statistic (mean, median) of pairwise distances, common in coreset and batch active learning (sener2018activelearningconvolutionalneural).
Uncertainty. Uncertainty indicates potential informativeness: high uncertainty suggests the model benefits from training on such samples. Typical measures include entropy of $M$’s predictive distribution or margin-based variants. We write the uncertainty functional as $\mathcal{U}$.
Loss. Relevance is captured by loss $M$ incurs on Bob’s data (yoo2019learning). If $y^{\prime}=M_{\theta}(x)$, then $\ell(y^{\prime},y)$ quantifies prediction-label discrepancy; larger loss indicates greater potential training value.
@@
For multi-class image classification, one can choose cross-entropy loss $\ell_{\mathrm{CE}}$, margin-based uncertainty $u=\frac{1}{k}\sum_{i}(\max p_{i}-\text{second-max }p_{i})$ (settles2009active), diversity $d=$ mean pairwise $\|x-x^{\prime}\|$ over $D_{R}^{x}$, and $f(l,u,d)=\alpha l+\beta u+\gamma d$ for weights $\alpha,\beta,\gamma$.
The exact choices of $\ell,\mathcal{U},\mathcal{D},f$ are decided case-by-case. For example, an untrained model may benefit from higher diversity, while a fine-tuned model may benefit from higher uncertainty or loss. The model owner tunes the scoring function to their needs; our protocol construction and security are oblivious to this choice.
4.2. Supplementary Components
Before presenting the full protocol, we briefly introduce three supplementary components of PrivaDE, shown in Fig. 1 preceding the two core components. These components form part of the optimization techniques of PrivaDE, which are described in greater detail in Sec. 5.
Model Distillation. The model owner may optionally perform distillation on their model $M$, training a smaller model that closely approximates the performance of the original. In the remainder of this section, we assume Alice’s model $M$ denotes this distilled (smaller) model. See Sec. 5.1 for more details.
Split Model. We require Alice, the model owner, to partition the model $M$ into three blocks: $M=C\circ B\circ A$, where block $B$ is revealed to Bob while $A$ and $C$ remain private to Alice. This design reduces cryptographic overhead while balancing the leakage of the protocol. The detailed security analysis and splitting methodology are presented in Sec. 5.2.
Representative Set Selection. We require Bob to select a representative subset $D_{R}\subset D_{B}$ before running the core protocol, specifically a $(d,\delta)$-representative subset:
Definition 4.1 (Representativeness of $D_{R}$ to $D_{B}$).
Let $D_{R}\subset D_{B}$, and fix a distance threshold $d>0$ and outlier tolerance $\delta\in[0,1]$. Then $D_{R}$ is $(d,\delta)$-representative if
$\Pr_{x\sim D_{B}}\!\bigl[\exists\,x^{\prime}\in D_{R}:\|x-x^{\prime}\|<d\bigr]\geq 1-\delta,$
where the probability is taken with respect to the empirical distribution over $D_{B}$.
We write $\mathcal{R}(D_{B})$ for any algorithm that, given $|D_{B}|=n>k$, returns $D_{R}\subset D_{B}$ of size $k$. Bob may use any appropriate algorithm $\mathcal{R}$; we only verify the representativeness of the selected subset.
To verify representativeness, we propose a challenge protocol $\mathsf{CP}$. Bob commits each $x_{i}\in D_{B}^{x}$ via $\mathtt{com}_{1},\dots,\mathtt{com}_{n}$ and publishes the representative set $D_{R}$ by revealing indices $I_{R}:=\{i|x_{i}\in D_{R}^{x}\}$. Alice samples random $i\in[n]$ and challenges Bob to prove in zero knowledge that committed $x_{i}$ is well represented by $D_{R}$. Bob recovers $x_{i}$, computes $d_{i}^{\prime}=\min_{x^{r}_{j}\in D_{R}}\|x^{r}_{j}-x_{i}\|$, and:
  • otherwise, replies $\mathsf{fail}_{i}$.
By repeating this challenge, Alice gains confidence that $D_{R}$ is representative of $D_{B}$. A detailed analysis of representative set selection and the challenge protocol is given in Sec. 5.3.
@@ 4.3. The PrivaDE Protocol
    Verified inference. The model is evaluated on $D_{R}$, and a corresponding zero-knowledge proof is used to verify computation.
@@
Formally, we prove that the protocol realizes $\mathcal{F}^{k}_{\mathsf{Score}}$, which securely computes Algorithm 1. In particular, Alice’s input (the model weights) and Bob’s input (dataset) are protected except for some intended leakage, to which the degree is primarily tuned by the parameter $k$ (size of the representative set). This leakage profile is fully specified by $\mathcal{F}^{k}_{\mathsf{Score}}$ in Fig. 2.
@@
Parameters: model $M=C\circ B\circ A$, loss $\ell$, uncertainty function $\mathcal{U}$, diversity function $\mathcal{D}$, scoring function $f$, representative set selection function $\mathcal{R}$, represenattive set size $k$, distance threshold $d$, outlier ratio $\delta$, commitment parameter $\mathsf{pp}$ (1) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives $I_{R}$ of size $k$ (indices of representative set) from $P_{2}$. (2) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives a list of challenge indices II from $P_{1}$, and sends $I$ to $P_{2}$. (3) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives $D^{x}_{B|I}$ (The set $D_{B}^{x}$ indiced by II) and $D^{x}_{R|I}$ (respective points in $D^{x}_{R}$ closest to each $D^{x}_{B|I}$) from $P_{2}$. For each $i\in I$ $(x_{i}\in D^{x}_{B|I})$, $\mathcal{F}^{k}_{\mathsf{Score}}$ checks that there exists some $x_{j}\in D_{R|I}^{x}$ such that $\|x_{i}-x_{j}\|<d$. If this check fails for at least $\delta\cdot|I|$ indices, the functionality sends $\mathsf{abort}$ to both $P_{1}$ and $P_{2}$. (4) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives $\theta_{A}$ from $P_{1}$ and $D^{x}_{R}$ (of size $k$) from $P_{2}$, then computes $\overrightarrow{a_{A}}\leftarrow A_{\theta_{A}}(D^{x}_{R})$ and sends $\overrightarrow{a_{A}}$ to $P_{2}$. (5) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives $\theta_{B}$ from $P_{1}$, sends $\theta_{B}$ to $P_{2}$, then computes $\overrightarrow{a_{B}}\leftarrow B_{\theta_{B}}(\overrightarrow{a_{A}})$ and sends $\overrightarrow{a_{B}}$ to $P_{1}$ and $P_{2}$. (6) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives $\theta_{C}$ from $P_{1}$, computes $\overrightarrow{y^{\prime}}\leftarrow C_{\theta_{C}}(\overrightarrow{a_{B}})$. (7) $\mathcal{F}^{k}_{\mathsf{Score}}$ receives $D^{x}_{B},D^{y}_{R}$ from $P_{2}$ and calculates the score: $l=\frac{1}{k}\sum_{i=1}^{k}\ell(y^{\prime}_{i},y_{i}),\quad u=\mathcal{U}(\{y^{\prime}_{1},\ldots,y^{\prime}_{k}\}),\quad d=\mathcal{D}(D_{B}^{x},D_{R}^{x})$, and finally $\phi\leftarrow f(l,u,d)$. It then sends $\phi$ to both $P_{1}$ and $P_{2}$.
Figure 2. Functionality $\mathcal{F}^{k}_{\mathsf{Score}}$ for computing the score of a dataset.
@@
We are now ready to formally describe the protocol PrivaDE that realizes $\mathcal{F}^{k}_{\mathsf{Score}}$ (Fig. 2). PrivaDE uses the following tools:
@@
  • $(\mathsf{SetupCom},\mathsf{Commit},\mathsf{Open})$ is a commitment scheme satisfying Definition 2.2.1.
  • $(\mathsf{SetupZKP},\mathcal{P},\mathcal{V})$ is an NIZKPoK system satisfying Definition 2.2.3, defined for the NP language $L_{\textsf{ML-Inf}}$. We consider two scenarios: (1) the model weights and outputs are hidden as part of the witness and the data point is public (blue); (2) the data point is hidden as part of the witness and the model weights and outputs are public (red). For each scenario, read the black text together with the text in the corresponding colour.
    Our proof-of-concept implementation uses EZKL (zkonduit_ezkl).
@@
  • $\Pi_{\mathsf{Inference}}$ is a 2-party protocol satisfying Definition 2.2.4, which securely realizes $\mathcal{F}_{\mathsf{Inference}}$ (Fig. 3) with malicious security. We instantiate this with SPDZ2k (Cramer2018SpdZ2k) in MP-SPDZ (mpspdz).
@@
The Setup phase. The setup phase is an offline stage before Alice and Bob run the protocol. In this phase:
@@
  • Alice and Bob communicate with a trusted third party to obtain cryptographic parameters $\mathsf{pp}\leftarrow\mathsf{SetupCom}(1^{\lambda})$ and $\mathtt{CRS}\leftarrow\mathsf{SetupZKP}(1^{\lambda})$.
@@
  • Bob verifies the authenticity of his dataset $D_{B}$ with a data authority, which issues a signature attesting to the authenticity of $D_{B}$.
@@
  • Alice may optionally perform model distillation locally, then interacts with a model authority that splits $M$ into $C\circ B\circ A$ and certifies that the model split is valid (Sec. 5.2.1).
  • Alice and Bob jointly fix the non-cryptographic public parameters listed in Table 1, decided out-of-band before the main protocol.
Table 1. Public protocol parameters agreed by Alice (model owner) and Bob (data owner) before running PrivaDE.
Symbol Description
$m_{\mathsf{CP}}$ Number of challenges in the CP protocol.
$d$ Distance threshold used in CP and in the $(d,\delta)$-representativeness condition.
$\delta$ Allowed outlier ratio in the representativeness guarantee.
$k$ Size of the representative set $D_{R}$ (i.e., $|D_{R}|=k$).
$R$ Public algorithm for selecting the representative set $D_{R}$ from $D_{B}$.
$\ell,\ \mathcal{U},\ \mathcal{D},\ f$ Public definition of the scoring function: loss $\ell$, uncertainty functional $\mathcal{U}$, diversity functional $\mathcal{D}$, and aggregator $f$.
@@
This is the only phase that involves a trusted third party. The main protocol then proceeds as in Fig. 5.
@@
Parties: Model owner $P_{1}$, data owner $P_{2}$ Private parameters: Dataset $D_{B}$ (held by $P_{2}$); model weights $\theta_{A},\theta_{C}$ (held by $P_{1}$) Public parameters: Commitment parameter $\mathsf{pp}$, representative set size $k$, distance threshold $d$, outlier ratio $\delta$, model architecture $M$, number of challenges $m_{\mathsf{CP}}$, model weights $\theta_{B}$ Stage 0: Commitments (1) $P_{1}$ computes $(\mathtt{com}_{A},r_{A})\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{A})$ and $(\mathtt{com}_{C},r_{C})\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{C})$, and sends $\mathtt{com}_{A},\mathtt{com}_{C}$ to $P_{2}$. (2) For each $(x_{i},y_{i})\in D_{B}$, $P_{2}$ computes $(\mathtt{com}_{x_{i}},r_{x_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},x_{i})$ and $(\mathtt{com}_{y_{i}},r_{y_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},y_{i})$, and sends $\mathtt{com}_{x_{i}},\mathtt{com}_{y_{i}}$ to $P_{1}$. Stage 1: Representative Set Selection (3) $P_{2}$ runs the representative set selection algorithm $\mathcal{R}$ locally to select a subset $D_{R}$ of size $k$ from $D_{B}$. It then computes $d^{\prime}$ as the $(1-\delta)$-th percentile of the pairwise distances $\Big\{\min_{b\in D_{R}^{x}}\|a-b\|:a\in D_{B}^{x}\setminus D_{R}^{x}\Big\}$. If $d^{\prime}>d$, $P_{2}$ aborts the protocol. Otherwise, he sends indices $I_{R}$ to $P_{1}$. (4) $P_{1}$ and $P_{2}$ runs the challenge protocol $\mathsf{CP}$, where $P_{1}$ supplies a list of challenge indices II of size m. If the challenge protocol fails, $P_{1}$ aborts. Stage 2: Secure Model Inference (Core Component 1) (5) $P_{1}$ and $P_{2}$ run the 2PC protocol $\Pi_{\mathsf{Inference}}$ to evaluate block AA on $D_{R}^{x}$. $P_{1}$ inputs $(\theta_{A},r_{A})$; $P_{2}$ inputs $D_{R}^{x}$ (and the corresponding opening randomness). $P_{2}$ receives the activations $\overrightarrow{a_{A}}=\{a_{A,i}\}_{i}$ and per-sample commitments $\{(\mathtt{com}_{a_{i}},r_{a_{i}})\}_{i}$; both parties obtain $\{\mathtt{com}_{a_{i}}\}_{i}$. (6) $P_{2}$ locally computes $\overrightarrow{a_{B}}\leftarrow B_{\theta_{B}}(\overrightarrow{a_{A}})$. He generates a ZK proof $\pi_{B}$ for the language $L_{\textsf{ML-Inf}}$ (public weights, hidden inputs) with witness $(\overrightarrow{a_{A}},\{r_{a_{i}}\}_{i})$, and sends $(\overrightarrow{a_{B}},\pi_{B})$ to $P_{1}$. (7) $P_{1}$ verifies $\pi_{B}$; if verification fails, she aborts. (8) $P_{1}$ computes $\overrightarrow{y^{\prime}}\leftarrow C_{\theta_{C}}(\overrightarrow{a_{B}})$ and for each $i$ computes $(\mathtt{com}_{y^{\prime}_{i}},r_{y^{\prime}_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},y^{\prime}_{i})$. She then produces a ZK proof $\pi_{C}$ for $L_{\textsf{ML-Inf}}$ (hidden weights, public inputs) with witness $(\theta_{C},r_{C},y^{\prime}_{1},\ldots,y^{\prime}_{k},r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}})$, and sends $\{\mathtt{com}_{y^{\prime}_{i}}\}_{i=1}^{k}$ and $\pi_{C}$ to $P_{2}$. (9) $P_{2}$ verifies $\pi_{C}$; if verification fails, he aborts. Stage 3: Dataset Scoring (Core Component 2) (10) $P_{1}$ and $P_{2}$ run the maliciously secure 2PC protocol $\Pi_{\mathsf{SubScore}}$. $P_{1}$ inputs $\big(y^{\prime}_{1},\ldots,y^{\prime}_{k},\,r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}},\,\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}\big)$; $P_{2}$ inputs $\big(D_{B}^{x},\,y_{1},\ldots,y_{k},\,r_{y_{1}},\ldots,r_{y_{k}},\,\mathtt{com}_{y_{1}},\ldots,\mathtt{com}_{y_{k}}\big)$. Both parties receive the score $\phi$.